# We assume the reader familiar with the definition of a partially ordered set, for which we will use the short form poset, and of the special case

### 19 Dec 2023 v1

@@ 1. Definitions and examples
Clearly, an embedding of posets is one-to-one, but a
@@
Indeed, if one chooses any total ordering of and the
opposite of that ordering, their intersection is
@@
A bijection from a -element antichain to the poset
is isotone and leaves the dimension, unchanged.
@@ 2. Our main result, and some consequences
The properties of vector-space dimension might
@@
the elements and agree in at least one
coordinate,
@@
still more general situation, that applies to the product
@@
actually be the dimension of but simply an integer such that
@@
for the original choices of clearly remains true for these
@@
since has a coordinate agrees with
@@
everywhere replaced by greatest elements and
@@
We need a bit more notation.
@@
Let us obtain one such result here as another consequence
@@ 3. An old result, and questions old and new
relative order under each of the orderings (3.2)
of the elements and
of
Reindexing those orderings if necessary, we may assume
without loss of generality that
for while for
We now claim that the map
@@
Since (3.4) is clearly an isotone map,
@@ C. Lin [11] poses part (iii) of that
question, though Theorems 10 and 11 of his paper
@@
And finally, a possible weaker implication than either of the above two:
@@
to get a positive answer to Question 3.2?
@@
Much more generally (but much more vaguely), let us ask
@@ 4. Absorbency
nor the corresponding condition with the roles of
@@
does equality in fact always hold the final inequality
of that statement?
@@
antichains, and I know of no cases not arising in a similar way.
@@
If the posets and are connected, must equality
hold in the final inequality of (4.5)?
@@
first part of the conclusion of (4.4)) would
@@ 5. Another function related to dimension, and a few from the literature
However, “bounded” is used in this way in the literature on
posets; e.g., [8, p.179, 12th line from bottom].)
@@
another function which can be useful in studying
the dimensions of posets.
@@
Thus, version (ii) of the definition of bounded dimension,
says that is the greatest of the
@@
and recalling that
@@
while the poset has dimension and absorbency
@@
Namely, in if on the one hand, precedes
@@
The term “Boolean dimension” refers to the fact that
can be expressed by a Boolean word
in variables.)

### 30 Dec 2023 v2

@@ 1. Definitions and examples
Clearly, an embedding of posets is one-to-one; but a
@@
Indeed, if one chooses any total ordering of then the
intersection of that ordering and the opposite ordering is
@@
On the other hand, a bijection from a -element antichain to
the poset is isotone and leaves the dimension, unchanged.
@@
In the verification, one takes a representation
for as in Definition 1.1(i), and
representations of the as in Definition 1.1(ii)
(which may, but do not need to satisfy the stronger condition (i)).
By repeating mappings if necessary, one can assume that the number of
chains used in each of these representations is the supremum
indicated above.
One can then easily combine these to get an embedding
of in a product of that number of chains.
@@ 2. Our main result, and some consequences
The properties of vector-space dimension
@@
the elements
agree in at least one coordinate,
@@
still more general situation, that applies to products
@@
be the dimension of (as it is in the
above example), but simply an integer such that
@@
for the original choices of remains true for these
@@
since has a coordinate must agree with
@@
everywhere replaced by greatest elements and ∎
@@
We will need a bit more notation.
@@
Let us obtain one such result here, as another consequence
@@ 3. An old result, and questions old and new
relative order, under each of the orderings (3.2),
of the elements and of
Reindexing those orderings if necessary, we may assume that
for some
for while
for
Let us now show that the map
@@
Since (3.4) is clearly isotone,
@@ C. Lin [11] poses part (iii) of that
question, though Theorems 10 and 11 of that paper
@@
And finally, a possible implication weaker than either of the above two:
@@
to get a positive answer to Question 3.2 for
finite posets?
@@
Much more generally (but much more vaguely), we may ask
@@ 4. Absorbency
nor the corresponding inequality with the roles of
@@
does equality always hold in the final inequality of that implication?
@@
antichains, and I know of no cases not arising in that way.
@@
If the posets and in (4.5)
are connected, must equality
hold in the final inequality of the conclusion thereof?
@@
first assertion of (4.4)) would
@@ 5. Another function related to dimension, and a few from the literature
However, “bounded” is used in this way in places in the literature;
e.g., [8, p.179, 12th line from bottom].)
@@
another function which is useful in studying dimensions.
@@
Thus, by version (ii) of the definition of bounded dimension,
is the greatest of the
@@
Then the left-hand side of (5.3)
is easily seen to be while the right-hand side is
@@
while the poset has dimension and, by (5.4),
absorbency
@@
Namely, in if and only
if on the one hand, precedes
@@
The term “Boolean dimension” refers to the fact that a map
can be expressed by a Boolean word
in variables.